# Leveraging MLLM Embeddings and Attribute Smoothing for Compositional Zero-Shot Learning

### v1

@@ 3.2.2 Attribute-Object Disentanglement
As mentioned before, one of the key challenges for CZSL task is to disentangle attribute and object from visual features. To overcome such challenge, we propose a novel weighted disentanglement module to disentangle primitives, as illustrated in Figure 2. For brevity, one image pair $x^{m}$ and $x^{a}$ from the triplet image set is taken as an example to elaborate on this module, while another image pair $x^{m}$ and $x^{o}$ follows the same architecture.
Weights computation. The features of $x^{m}$ and $x^{a}$ (i.e., $F^{a}$ and $F^{o}$) are vertically concatenated and fed into two MLP modules to derive their respective weights of shared attribute features relative to each other, and subsequently utilize them to compute the weights of their own exclusive object features as follows:
@@
where $w_{attr}^{m2a},w_{attr}^{a2m}\in\mathbb{R}^{h}$ demonstrate the weights of the shared attribute features of $x^{m}$ relative to $x^{a}$, and $x^{a}$ relative to $x^{m}$, respectively. $w_{obj}^{m2a}$ and $w_{obj}^{a2m}$ denote the weights of exclusive object features corresponding to $x^{m}$ and $x^{a}$, respectively, which are derived by ”$1-shared\ weights$” paradigm as beyond the shared features of the image pair are the exclusive features of each image. Taking $w_{attr}^{m2a}$ as an example, its $k$-th element refers to the shared attribute proportion of $k$-th feature of $x^{m}$ relative to $x^{a}$.
Disentangled features obtainment. We multiply elements of each weight by the corresponding features and then calculate the average. The following takes the process of obtaining the shared attribute features of image $x^{m}$ relative to $x^{a}$ as an example:
$f_{attr}^{m2a}=\frac{1}{h}\sum_{i=1}^{h}{w^{m2a}_{attr}}_{i}\ {F^{a}}_{i,:}$ (6)
where ${F^{a}}_{i,:}$ denotes the $i$-th row of $F^{a}$, i.e., the $i$-th feature of $x^{a}$. ${w^{m2a}_{attr}}_{i}$ refers to the $i$-th element of $w^{m2a}_{attr}$, and $f_{attr}^{m2a}\in\mathbb{R}^{d}$ is the shared attribute feature of $x^{m}$ relative to $x^{a}$.
For the image pair of $x^{m}$ and $x^{a}$, four parts are obtained: the shared attribute features of $x^{m}$ relative to $x^{a}$, and $x^{a}$ relative to $x^{m}$, as well as two exclusive object features of the image pair, respectively. These four features are marked as $f_{pri}^{e}$, where $e\in\{m2a,a2m\}$ and $pri\in\{attr,obj\}$. Then the shared attribute feature of $x^{a}$ and $x^{m}$ without relativity is obtained by an MLP layer, which is less dependent on the object. The process is as follows:
@@
Similarly, we disentangle attribute and object for $x^{m}$ and $x^{o}$ and obtain the same features as $x^{m}$ and $x^{a}$: $f_{pri}^{e}$, where $e\in\{m2o,o2m\}$ and $pri\in\{obj,attr\}$ and $f_{obj}^{mo}$.
@@ 3.2.3 Feature Alignment
Inspired by [24] that leverages the last hidden states as the representation embeddings, we consider the last hidden states of LLaVA v1.5 [16] as our MLLM embeddings for words. Moreover, to tackle the problem that the ineffective overconfidence exhibited by the models in terms of the ground-truth attribute hinders them from generalizing to unseen compositions, GPT 3.5 is employed to generate several auxiliary attributes that describe an object with only one ground-truth attribute and perform label smoothing during attribute alignment. Now we detail each part of feature alignment.
Generating auxiliary attribute words by LLM. Since only attribute text needs to be generated, we leverage a LLM, GPT-3.5, instead of MLLM, to generate several auxiliary attributes for each composition. Specifically, the following prompt is input to LLM: ’Please give me $t$ adjectives that can describe the visual feature of a photo of a/an … well.’, where $t$ is the number of auxiliary attributes and attribute-object composition (e.g., peeled apple) is filled in ’’. Please refer to Appendix A for more details about the generation of auxiliary attributes by GPT-3.5. Subsequently, the generated auxiliary attribute words form a set $\mathcal{A}_{a}$. Therefore, the set of all words $\mathcal{Y}$ is obtained, including attributes, objects and auxiliary attributes as follows:
@@
Obtaining MLLM embeddings for words and compositions. Each word $y\in\mathcal{Y}$ is fed into LLaVA v1.5 to get the last hidden states, i.e., $LLaVA_{lhs}(\cdot)$. Please refer to Appendix B for more details about the obtainment of the last hidden states of LLaVA v1.5 for an input word. Subsequently, they are passed through an MLP layer to get embeddings $E_{word}(\cdot)$ of aligned dimension with visual features. And for a composed pair $c$ of attribute $a$ and object $o$, i.e., $c=(a,o)$, we get the last hidden states of LLaVA v1.5 for $a$ and $o$, respectively, which are then horizontally concatenated and fed into a linear layer $Lin_{co}(\cdot)$ to get the composed pair embedding $E_{co}(\cdot)$. The process is formulated as follows:
$E_{word}(y)=MLP_{wd}(LLaVA_{lhs}(y))$ (9)
$E_{co}(c)=Lin_{co}(Cat((LLaVA_{lhs}(a),(LLaVA_{lhs}(o)))$ (10)
Word expanding. Prior works compute cosine similarities of disentangled features and word embeddings only within the respective domains of attributes or objects, which results in the disentangled attributes and objects still retaining the information of each other. To address the problem, we propose a word expanding strategy, which computes cosine similarities of visual features and the embeddings of all words, including attributes and objects, and treats all words except the ground-truth word as negative labels.
Alignment by cross-entropy. Similar to [19], we use cross-entropy to process the cosine similarity of visual features and word embeddings. Assume that $f$ is the visual embedding and $E_{word}(wd)$ is the word embedding for the word $wd\in\mathcal{Y}$ in a joint space. The classifier logit from $f$ to $E_{word}(wd)$ is defined as follows:
$CE(f,wd)=\frac{e^{\delta\cdot cos(f,E_{word}(wd))}}{\sum_{y\in\mathcal{Y}}e^{\delta\cdot cos(f,E_{word}(y))}}$ (11)
where $\delta$ is the temperature variable, and $cos(\cdot,\cdot)$ denotes cosine similarity function. Thus cross-entropy with/without label smoothing can be uniformly formulated as follows:
@@
where $\alpha$ denotes the smoothing factor, $t$ refers to the number of auxiliary labels and $z\in[0,1]$ represents the target value of one-hot or smoothing label. For cross-entropy without label smoothing, i.e. with one-hot label $H_{oh}$, $\alpha$ is set to 00. And the cross-entropy with label smoothing is denoted as $H_{ls}$.
For the disentangled attribute features of one image relative to each other, since a single object exhibits multiple attributes, we exploit attribute smoothing with auxiliary attributes to undermine the confidence in the ground-truth attribute and learn more related attributes. For the shared attribute features without relativity, one-hot label is used to compute its classification loss. The loss for disentangled attributes can be defined as follows:
@@
Concerning the disentangled object features, we use cross-entropy with one-hot label to learn the prototype of the object and the loss is as follows:
@@
With respect to the visual feature of the image from image embedder, we calculate the cosine similarity between visual embedding and the composed pair embedding of the corresponding composition label and use one-hot label to align them. The classification loss for compositions is as follows:
$\mathcal{L}_{comp}=\sum_{img\in\{m,a,o\}}H_{oh}(F_{comp}^{img},\mathcal{C}_{s})$ (15)
@@ 3.3 Training and Inference
During the training phase, the overall loss function is formulated as follows:
$\mathcal{L}=\gamma_{ortho}\mathcal{L}_{ortho}+\gamma_{comp}\mathcal{L}_{comp}+\gamma_{attr}\mathcal{L}_{attr}+\gamma_{obj}\mathcal{L}_{obj}$ (16)
where $\gamma_{ortho}$, $\gamma_{comp}$, $\gamma_{attr}$, and $\gamma_{obj}$ are weighting factors to balance the influence of different losses.
For inference, we use the composition feature space generated by the classifier that is obtained by optimizing $\mathcal{L}_{comp}$. Specifically, given an image from test set, the cosine similarities of its visual feature obtained by image embedder and the composed pair embeddings of all candidate compositions in the test set are computed. The composition with the highest similarity is the class predicted by the model. Note that although the disentanglement branches are not used for inference, they still influence the formation of the composition feature space through their shared visual feature extraction module described in Section 3.2.1.
MIT-States C-GQA VAW-CZSL
Method $AUC$ $HM$ $Seen$ $Unseen$ $AUC$ $HM$ $Seen$ $Unseen$ $AUC$ $HM$ $Seen$ $Unseen$
SymNet [15] 3.2 13.7 22.7 20.1 1.9 10.8 20.3 11.8 2.8 13.5 20.2 18.0
CompCos [19] 12.3 28.2 39.0 39.5 5.0 17.7 32.8 19.1 6.5 20.8 30.5 27.4
Co-CGE [20] 10.3 25.1 41.0 33.1 4.2 15.2 32.9 17.0 6.2 19.7 31.0 26.1
SCEN [13] 9.8 24.6 35.1 36.5 3.8 15.3 31.5 15.7 5.7 19.2 29.9 24.5
OADis [38] 13.1 29.0 42.3 27.3 2.3 12.1 23.3 12.8 4.1 16.2 26.0 20.7
INV [44] 11.5 26.6 28.5 25.0 1.4 7.9 28.6 6.8 2.0 11.1 21.1 11.9
CANet [42] 13.6 29.8 46.4 39.9 5.7 18.9 34.8 20.5 6.7 21.0 31.2 27.4
ProCC [10] 9.5 28.1 43.1 39.1 3.5 15.1 32.4 15.8 3.6 18.9 26.9 25.5
CLIP [28] 11.0 26.1 30.2 46.0 1.4 8.6 7.5 25.0 - - - -
CoOp [28] 13.5 29.8 34.4 47.6 4.4 17.1 20.5 26.8 - - - -
@@
Table 1: Comparison with the state-of-the-art results on MIT-States, C-GQA and VAW-CZSL. We compare our TRIDENT with the state-of-the-art methods on test $AUC$, best seen ($Seen$), best unseen ($Unseen$) and best harmonic mean ($HM$) accuracies on these three datasets. We measure top-1 $AUC$ on MIT-States and C-GQA, and top-3 $AUC$ on VAW-CZSL. Best results are displayed in boldface, and second best results are underlined.
4 Experiment
4.1 Experiment Setup
Datasets. We evaluate our model on three challenging CZSL benchmark datasets: MIT-states [11], C-GQA [25], and VAW-CZSL [38]. We present the introduction and common data splits of the three datasets in Appendix C.
Metrics. Following the common generalized CZSL setting [33], we evaluate our model on seen and unseen pairs separately. Based on them, a calibration bias trades off between the accuracies of seen and unseen pairs. We calculate area under the curve $AUC$ (in %) using seen and unseen classification accuracies at different biases in test data. The best seen and unseen accuracies $Seen$ and $Unseen$ (in %) of the curve are also reported. In addition, we calculate the harmonic mean of seen and unseen classification accuracies at difference biases and report the best one $HM$ (in %).
Implementation details. We use the visual encoder of LLaVA v1.5, Vit-large-14-336px as our frozen feature extractor, whose outputs contain 577 tokens (1 [CLS] and 576 patch tokens) of 1024 dimensions. The cross-modal connector of LLaVA v1.5 maps the features to the dimension of 4096, the same as last hidden states of based LLM Vicuna v1.5 [45]. Image embedder and the MLP for words map them to the dimension of 1024 for faster training. TRIDENT and all baseline models are trained with 128 batch size for 50 epochs. The number of global features is set to 6, 2, 4 for the three datasets, respectively, and the number of local features is twice that of global features. The label smoothing factor is set to 0.09, 0.03, 0.03 for the three datasets, respectively. The number of generated auxiliary attributes for each composition is set to 3. Refer to Appendix D for more information about implementation.
Baselines. We compare our TRIDENT with recent and prominent approaches in the task of CZSL: SymNet [15], CompCos [19], Co-CGE [20], SCEN [13], OADis [38], INV [44], CANet [42], and ProCC [10]. We replace their backbone with Vit-large-14-336px and retrain all models with the same epoch for the sake of fairness. In addition, although comparing TRIDENT with CLIP-based CZSL methods, which rely on the dual-tower architecture, is very unfair due to the significant addition of trainable parameters and training overhead for both the text and visual encoders, we still choose the foundational CLIP and CoOp models as baselines for their strong zero-shot classification abilities.
4.2 Results and Discussion
In this section, we compare TRIDENT with state-of-the-art methods. As shown in Table 1, TRIDENT surpasses other models by a substantial margin in general. For MIT-States, TRIDENT boosts $AUC$, $HM$, and $Unseen$ from 13.6%, 29.8%, and 39.9% of CANet to the new state-of-the-art performance of 14.2%, 30.9%, and 40.0% with 0.6%, 1.1%, and 0.1% improvement, respectively. Our model achieves competitive performance on MIT-States benchmark, despite considerable label noise [2]. However, for the more challenging benchmark C-GQA, TRIDENT achieves 8.0%, 22.6%, 39.5%, and 24.1% on the metrics of $AUC$, $HM$, $Seen$, and $Unseen$, providing 2.3%, 3.7%, 4.7%, and 3.6% improvements on the previous state-of-the-art model CANet. For the existing most challenging benchmark dataset VAW-CZSL, TRIDENT attains performance of 8.3%, 23.4%, 23.4%, and 33.3%, surpassing CANet by 1.6%, 2.4%, 2.2%, and 3.7% in terms of $AUC$, $HM$, $Seen$, and $Unseen$. The largest improvement is observed in the $Unseen$ metric, indicating that attribute smoothing helps enhance the generalization ability of the model. We observe TRIDENT performs significantly better than CANet regarding all metrics on two challenging and low-noise benchmark dataset C-GQA and VAW-CZSL, indicating the efficacy of our approach. This improvement arises from the utilization of MLLM embeddings and attribute smoothing, which enhance attribute-object disentanglement and consequently facilitate the recognition of unseen compositions while maintaining performance on seen compositions.
In addition, we compare TRIDENT with dual-tower CLIP and CoOp after fine-tuned for the CZSL task. Since they are trained on a large amount of image-text data, they possess zero-shot image classification capabilities, which leads to better classification results for unseen images. Regarding $Unseen$ metric, CoOp outperforms TRIDENT by 7.6% and 2.7% on MIT-States and C-GQA, respectively. However, TRIDENT surpasses CoOp by 0.7% and 1.1% on the core metrics of $AUC$ and $HM$ on MIT-States, as well as 3.6% and 5.5% on C-GQA, which suggests TRIDENT performs better than CLIP and CoOp in CZSL task.
4.3 Ablation Study
@@
(a) image-to-text retrieval.
(b) text-to-image retrieval (successful cases).
(c) text-to-image retrieval (failure cases).
Figure 3: Qualitative analysis. (a) Top-5 image-to-text retrieval cases. The first two rows display successful cases, while the last row presents failure cases. (b) Successful cases of top-5 text-to-image retrieval. (c) Failure cases of top-5 text-to-image retrieval. In all cases, the attribute and object of composition label are marked in orange and blue, respectively. And the successful and failure retrieval results are tagged in green and red, respectively.
Effectiveness of each component. We ablate certain module of TRIDENT to evaluate the contribution of each module on MIT-States, as it is the most common used dataset. The ablation results are reported in Table 2. From this table, we gain the following observations.
1) Both w/o condition_masks model and w/o FAAs model perform worse than TRIDENT, which validates the importance of extracting the multi-granularity features and filtering out the background noise, respectively.
2) TRIDENT surpasses w/o word_expanding model and w/o attribute_smoothing model on the $Unseen$ metric, yet falls short of them on the $Seen$ metric. The difference between TRIDENT and the w/o word_expanding model stems from its more thorough disentanglement, which enhances the recognition of unseen images while weakens the identification of seen images. The disparity between TRIDENT and the w/o attribute_smoothing model arises from attribute smoothing, which diminishes the confidence of the model in seen compositions, facilitating its generalization to unseen compositions. However, the improvement of TRIDENT over these two models on $AUC$ and $HM$ indicates the effectiveness of word expanding and label smoothing strategy.
3) TRIDENT outperforms w/o $\mathcal{L}_{attr}+\mathcal{L}_{obj}$ model on all metrics, confirming that the attribute-object disentanglement module is highly advantageous for generalization from seen compositions to unseen compositions.
4) w/o $\mathcal{L}_{ortho}$ model is inferior to TRIDENT, which suggests the designed orthogonal regularization is helpful to guarantee different features extract different information.

### v2

@@ 3.2.2 Attribute-Object Disentanglement
As mentioned before, one of the key challenges for CZSL is to disentangle attributes and objects from visual features. To overcome the challenge, we propose a novel weighted disentanglement strategy, as illustrated in Figure 2. For brevity, the image pair $(x^{m},x^{a})$ from the triplet image set is taken as an example to elaborate on this strategy, while another image pair $(x^{m},x^{o})$ is processed in the same manner.
Weights computation. The features of $x^{m}$ and $x^{a}$ (i.e., $F^{m}$ and $F^{a}$) are vertically concatenated and fed into two MLP modules to derive their respective weights of shared attribute features relative to each other. Subsequently, we utilize them to compute the weights of their own exclusive object features as follows:
@@
where $w_{attr}^{m2a},w_{attr}^{a2m}\in\mathbb{R}^{h}$ represent the weights of the shared attribute features of $x^{m}$ relative to $x^{a}$ and $x^{a}$ relative to $x^{m}$, respectively. $w_{obj}^{m2a}$ and $w_{obj}^{a2m}$ denote the weights of exclusive object features corresponding to $x^{m}$ and $x^{a}$, respectively, which are derived by ”$1-shared\ weights$” paradigm as beyond the shared features of the image pair are the exclusive features of each image. Taking $w_{attr}^{m2a}$ as an example, its $k$-th element refers to the shared attribute proportion of the $k$-th feature of $x^{m}$ relative to $x^{a}$.
Disentangled features acquisition. We multiply the elements of each weight vector by the corresponding features and then calculate the average. The following takes the process of obtaining the shared attribute features of image $x^{m}$ relative to $x^{a}$ as an example:
$f_{attr}^{m2a}=\frac{1}{h}\sum_{i=1}^{h}{w^{m2a}_{attr}}_{i}\ {F^{m}}_{i,:}$ (6)
where ${F^{m}}_{i,:}$ denotes the $i$-th row of $F^{m}$, i.e., the $i$-th feature of $x^{m}$. ${w^{m2a}_{attr}}_{i}$ refers to the $i$-th element of $w^{m2a}_{attr}$.
For the image pair of $x^{m}$ and $x^{a}$, four parts are obtained: the shared attribute features of $x^{m}$ relative to $x^{a}$ and $x^{a}$ relative to $x^{m}$, as well as two exclusive object features of the two images, respectively. These four features are marked as $f_{pri}^{e}$, where $e\in\{m2a,a2m\}$ and $pri\in\{attr,obj\}$. Then the shared attribute feature of $x^{a}$ and $x^{m}$ without relativity is obtained by an MLP layer, which is less dependent on the different objects of the two images. The process is as follows:
@@
Similarly, we disentangle the attribute and object for the image pair ($x^{m}$, $x^{o}$) and obtain the same visual features as ($x^{m}$, $x^{a}$): $f_{pri}^{e}$, where $e\in\{m2o,o2m\}$ and $pri\in\{obj,attr\}$, and the feature without relativity $f_{obj}^{mo}$.
@@ 3.2.3 Feature Alignment
Inspired by GritLM Muennighoff et al. (2024) that leverages the last hidden states of LLMs as the representational embeddings, we consider the last hidden states of LLaVA v1.5 Liu et al. (2024a) as our MLLM embeddings for primitive words. Moreover, to tackle the problem that the ineffective overconfidence exhibited by the model in terms of the ground-truth attribute hinders it from generalizing to unseen compositions, we employ GPT-3.5 OpenAI (2023) to generate auxiliary attributes for compositions and perform label smoothing during attribute alignment. The auxiliary attributes and MLLM embeddings are obtained offline before training TRIDENT.
Auxiliary attributes generation by LLM. Since only textual attributes need to be generated, the LLM GPT-3.5 OpenAI (2023), instead of an MLLM, is leveraged to generate $t$ auxiliary attributes for each composition. Specifically, the following prompt is input to LLM: ’Please give me $t$ adjectives that can describe the visual feature of a photo of a/an … well.’, where the attribute-object composition (e.g., peeled apple) is filled in ’’. Subsequently, the generated auxiliary attribute words form a set $\mathcal{A}_{a}$. Therefore, the set of all words $\mathcal{Y}$ is obtained, including attributes, objects and auxiliary attributes:
@@
MLLM embeddings acquisition. Each word $y\in\mathcal{Y}$ is fed into LLaVA v1.5 to get the last hidden states, i.e., $LLaVA_{lhs}(\cdot)$. Specifically, $y$ is tokenized into multiple sub-words and passed through the MLLM; the final-layer output is averaged as the MLLM embedding of the word. Subsequently, it is passed through an MLP layer to obtain the embedding $E_{word}(\cdot)$ of the aligned dimension with visual features. And for a composed pair $c$ of attribute $a$ and object $o$, i.e., $c=(a,o)$, we horizontally concatenate the MLLM embeddings for $a$ and $o$ and feed them into a linear layer $Lin_{co}(\cdot)$ to get the composed pair embedding $E_{co}(\cdot)$. The process is formulated as follows:
$E_{word}(y)=MLP_{word}(LLaVA_{lhs}(y))$ (9)
$E_{co}(c)=Lin_{co}(Cat(LLaVA_{lhs}(a),(LLaVA_{lhs}(o)))$ (10)
Word expanding. Previous works compute cosine similarities of disentangled features and word embeddings and apply cross-entropy only within the respective domains of attributes or objects, which results in the disentangled attributes and objects still retaining the information of each other. To address the problem, we propose a novel word expanding strategy, which computes cosine similarities of visual features and the embeddings of all words, including attributes and objects, and treats all words except the ground-truth word as the negative labels in subsequent cross-entropy.
Alignment by cross-entropy. Similar to Mancini et al. (2021), we use cross-entropy to align the visual features and word embeddings. Assume that $f$ is the visual feature and $E_{word}(wd)$ is the word embedding for the word $wd\in\mathcal{Y}$. The classifier logit from $f$ to $E_{word}(wd)$ is defined as follows:
$CE(f,wd)=\frac{e^{\delta\cdot\cos(f,E_{word}(wd))}}{\sum_{wd^{{}^{\prime}}\in\mathcal{Y}}e^{\delta\cdot\cos(f,E_{word}(wd^{{}^{\prime}}))}}$ (11)
where $\delta$ is the temperature variable, and $cos(\cdot,\cdot)$ denotes the cosine similarity function. Thus cross-entropy with/without label smoothing can be uniformly formulated as follows:
@@
where $\alpha$ denotes the smoothing factor, $t$ refers to the number of auxiliary labels, and $z\in[0,1]$ represents the target value of one-hot label or smoothing label. For cross-entropy without label smoothing, i.e., with one-hot label $H_{oh}$, $\alpha$ is set to 00. The cross-entropy with label smoothing is denoted as $H_{ls}$.
For the disentangled attribute features of one image relative to each other, since a single object exhibits multiple attributes, we exploit attribute smoothing with auxiliary attributes to undermine the overconfidence in the ground-truth attribute and learn more related attributes. For the shared attribute features without relativity, the one-hot label is used to learn a pure attribute concept that is less conditioned on objects. The loss for disentangled attributes is as follows:
@@
Concerning the disentangled object features, we use cross-entropy with one-hot label to learn the object concept and the loss is as follows:
@@
With respect to the visual feature of the composition from image embedder, we calculate the cosine similarities between the visual embedding and the composed pair embeddings of seen composition labels. The cross-entropy loss for the composition is as follows:
$\mathcal{L}_{comp}=\frac{1}{|img|}\sum_{img\in\{m,a,o\}}H_{oh}(f_{comp}^{img},\mathcal{C}_{s})$ (15)
@@ 3.3 Training and Inference
During the training phase, the overall loss is as follows:
$\mathcal{L}=\gamma_{ortho}\mathcal{L}_{ortho}+\gamma_{comp}\mathcal{L}_{comp}+\gamma_{pri}(\mathcal{L}_{attr}+\mathcal{L}_{obj})/2$ (16)
where $\gamma_{ortho}$, $\gamma_{comp}$, and $\gamma_{pri}$ are weighting factors to balance the influence of different losses.
For inference, given an image from the test set, the cosine similarities of its visual feature obtained by image embedder and the composed pair embeddings of all candidate compositions are computed. The composition with the highest similarity is predicted by the model. Note that although the disentanglement branch is not used for inference, it still influences the formation of the composition feature space through the visual feature extraction module described in Section 3.2.1.
Dataset Primitive Train Validation Test
$|\mathcal{A}|$ $|\mathcal{O}|$ $|\mathcal{C}_{s}|$ $|\mathcal{X}|$ $|\mathcal{C}_{s}|$ $|\mathcal{C}_{u}|$ $|\mathcal{X}|$ $|\mathcal{C}_{s}|$ $|\mathcal{C}_{u}|$ $|\mathcal{X}|$
MIT-States 115 245 1262 30k 300 300 10k 400 400 13k
C-GQA 413 674 5592 27k 1252 1040 7k 888 923 5k
VAW-CZSL 440 541 1252 72k 2121 2322 10k 2449 2470 11k
Table 1: Summary statistics of the datasets used in our experiments.
Method MIT-States C-GQA VAW-CZSL
$AUC$ $HM$ $Seen$ $Unseen$ $AUC$ $HM$ $Seen$ $Unseen$ $AUC$ $HM$ $Seen$ $Unseen$
SymNet Li et al. (2020) 3.2 13.7 22.7 20.1 1.9 10.8 20.3 11.8 2.8 13.5 20.2 18.0
CompCos Mancini et al. (2021) 12.3 28.2 39.0 39.5 5.0 17.7 32.8 19.1 6.5 20.8 30.5 27.4
Co-CGE Mancini et al. (2022) 10.3 25.1 41.0 33.1 4.2 15.2 32.9 17.0 6.2 19.7 31.0 26.1
SCEN Li et al. (2022) 9.8 24.6 35.1 36.5 3.8 15.3 31.5 15.7 5.7 19.2 29.9 24.5
OADis Saini et al. (2022) 13.1 29.0 42.3 27.3 2.3 12.1 23.3 12.8 4.1 16.2 26.0 20.7
INV Zhang et al. (2022) 11.5 26.6 28.5 25.0 1.4 7.9 28.6 6.8 2.0 11.1 21.1 11.9
CANet Wang et al. (2023) 13.6 29.8 46.4 39.9 5.7 18.9 34.8 20.5 6.7 21.0 31.2 27.4
ProCC Huo et al. (2024) 9.5 28.1 43.1 39.1 3.5 15.1 32.4 15.8 3.6 18.9 26.9 25.5
CLIP Nayak et al. (2023) 11.0 26.1 30.2 46.0 1.4 8.6 7.5 25.0 - - - -
CoOp Nayak et al. (2023) 13.5 29.8 34.4 47.6 4.4 17.1 20.5 26.8 - - - -
@@
Table 2: Comparison with the state-of-the-art results on MIT-States, C-GQA and VAW-CZSL. The four indicators are explained in Metrics. We measure top-1 $AUC$ on MIT-States and C-GQA, and top-3 $AUC$ on VAW-CZSL. The best results are displayed in boldface, and the second best results are underlined.
@@
4 Experiment
4.1 Experiment Setup
Datasets. We evaluate our model on three challenging CZSL datasets: MIT-states Isola et al. (2015), C-GQA Naeem et al. (2021), and VAW-CZSL Saini et al. (2022). The common data splits are presented in Table 1.
Metrics. Our method is evaluated on seen and unseen compositions separately. Following the common generalized CZSL setting Purushwalkam et al. (2019), a calibration bias trades off between the accuracies of seen and unseen compositions. We calculate the Area Under the Curve ($AUC$) using seen and unseen accuracies at different biases. The best $Seen$ and $Unseen$ accuracies of the curve are also reported. In addition, we calculate the Harmonic Mean of seen and unseen accuracies at different biases and report the best one ($HM$).
Implementation details. We use the visual encoder of LLaVA v1.5, ViT-Large-Patch14-336px as our frozen visual backbone. TRIDENT and all baseline models are trained with the batch size of 128 for 50 epochs under the PyTorch framework Paszke et al. (2019). The number of global features is set to 6, 2, and 4 for the three datasets, respectively, and the number of local features is twice that of the global features. The label smoothing factor is set to 0.09, 0.03, and 0.03 for the three datasets, respectively. The number of auxiliary attributes generated for each composition is set to 3. We train TRIDENT by Adam optimizer with the weight decay of 5e-5, learning rates of 1.5e-6 for word embedding and 2e-4 for other modules. We decay the learning rate by a factor of 0.1 at epoch 30 and 40. The temperature variable of cosine similarity $\delta$ is set to 0.05. For weighting coefficients $\gamma_{ortho}$, $\gamma_{comp}$, and $\gamma_{pri}$, we set them to 0.1, 1, 0.25, respectively.